# TTD: Text-Tag Self-Distillation Enhancing Image-Text Alignment in CLIP to Alleviate Single Tag Bias

### v1

@@ 4.1.1 Multi-Tag Selection
While image-tag scoring ($s_{\mathrm{image}}$) demonstrates high precision, it experiences very low recall due to the single tag bias, often leading to the selection of only one tag so if misses other image-related tags. Similarly, text-tag scoring ($s_{\mathrm{text}}$) and their combination show comparable issues. However, our approach significantly mitigates the single tag bias, showing high recall. Also, it exhibits high accuracy and mAP, which indicates the effectiveness of our scoring method. We further examine sensitivity to different thresholds and alternative scoring techniques that incorporate external data beyond the image-text pair in Sec. 0.B.1.
@@ 4.2 Results on Open-Vocabulary Semantic Segmentation
Tab. 4 contrasts the tag-level segmentation performance of studies employing only image-text pairs without additional annotations or external models akin to ours for fair comparison. For this reason, models [23, 11, 27, 13] that utilize additional annotations or external models (e.g. mask proposal networks) are also excluded. Additionally, we compared models that utilize NLP-based $\mathcal{L}$ in the table) for tag extraction. Implementing our fine-tuning technique results in an average increase in mIoUs of 8.5% for MaskCLIP and 3.5% for TCL. Furthermore, TCL enhanced with our method outperforms existing models on benchmarks such as Pascal VOC, Pascal Context, and COCO-Object. This indicates that our approach successfully augments model alignment in addition to alleviating single tag bias.
@@ 4.3 Ablation
Table 5: Ablation studies
(a) Effect of loss terms
@@
(b) Effect of tagging method
@@ 4.3.2 Effect of Tag Selection Method
In our approach, tag selection is conducted via pixel-tag scoring (Eq. 2) to extract pseudo tags representing the image-text relationship. Alternatively, we examine NLTK and simple image-tag scoring (Eq. 1) for tag selection, as shown in Tab. 5(b). When employing NLTK to extract nouns from texts, due to the aforementioned issues (extraction of image-irrelevant tags and omission of image-relevant tags), the model struggles to generate the ideal text-image similarity map (i.e., the union of pseudo-tag driven maps). Consequently, the generated union map may include the activation of unrelated tags or omit activations necessary for the image-tag relationship, resulting in minor improvement. On the other hand, using image-tag scoring tends to extract a single tag, exacerbating single tag bias and resulting in a decline in both CaptionIoU and mIoU. In contrast, our tag selection method exhibits significantly enhanced performance in both tasks, indicating that the pseudo tags extracted using our method effectively capture the relationship between image and text.
@@ 5 Conclusion
Our research addresses the misalignment between images and text inherent in existing CLIP-based models. Specifically, we present a fine-tuning technique that resolves the issue where image and text representations are overly focused on a single tag, thereby enhancing both image segmentation and multi-tag classification within the text. This was achieved using only image and text pairs, without additional annotations. The key outcome of addressing the single tag bias is enabling CLIP to correctly identify all truth tags related to the image within the text, akin to how class activation map studies provide pixel-level class information without direct input to the network. Our TTD training method, which does not employ external modules for tag extraction, ensures that the predicted tags by the fine-tuned model have direct relevance to the image. This feature can effectively address the hallucination issue seen in conventional tag generation models, where natural language processing-based models might generate tags unrelated to the target image. Hence, our approach is valuable as it offers a solution to the hallucination problem, suggesting significant potential for various applications.
Appendix 0.A Experimental Details
0.A.1 Implementation Details
@@
We conduct model training using 4 NVIDIA A100 GPUs, each having 80GB of memory. Implementation details are summarized in Tab. 6. Most settings, including image size, prompt template usage, and post-processing, align with the original setup of baseline models [44, 4]. Our batch size is set to 32, the maximum feasible size given our available GPU resources. Training data consists of the CC3M and CC12M datasets [34, 5], with additional sample pruning to remove noisy samples. We calculate cosine similarity between image and text embeddings for all pairs, sampling those with similarity values exceeding the mean plus the standard deviation. Ultimately, 1.8M pairs are selected from a total of 12.8M pairs, as shown in Tab. 7. Training time has been reduced approximately sevenfold, with a performance boost of about +0.6 mIoU, underscoring the efficacy of filtering out noisy samples.
@@ 0.A.2 Evaluation Details
Appendix 0.B Additional Experiments
@@
0.B.1 Effect of tagging methods
@@
0.B.2 Multi-Tag Selection
Here, we delve into an additional scoring method not covered in the main paper. We explore the performance of previously introduced scoring techniques ($s_{\mathrm{image}}$ (Eq. 1), $s_{\mathrm{text}}$ (Eq. 7)) across various thresholds in Sec. 0.A.2.
As mentioned in Sec. 1, each tag can be scored based on the proportion of image pixels similar to that tag ($s_{\mathrm{seg}}$). Specifically, as depicted in Fig. 9, we generate a similarity map between the image and each tag. Subsequently, for every pixel, we identify the most relevant tag with the highest similarity value, enabling us to score each tag based on the proportion of its relevant pixels. As illustrated in Tab. 8, $s_{\mathrm{seg}}$ demonstrates relatively high precision since it accentuates a single tag. However, due to the tendency of irrelevant tags to occupy more pixels than relevant ones, it encounters challenges in score ranking, resulting in a significantly low mAP (i.e., 57.9). In contrast, our scoring method ($s_{\mathrm{pixel}}$) effectively mitigates the single tag bias while maintaining correct ranking, resulting in high F1 score and mAP.
@@ 0.B.2 Multi-Tag Selection
ADDS [40] O ViT-B/32 36.56
MKT [15] O ViT-B/16 37.6
DualCoOp [36] O ResNet-50 43.6
TaI-DPT [14] O ResNet-50 44.99
CLIP [31] X ResNet-50 35.58
CLIP Surgery [22] X ResNet-50 39.99
CLIP Surgery [22] X ViT-B/16 40.53
Ours (TCL [4]+TTD) X ViT-B/16 42.63
0.B.3 Open-Vocabulary Multi-label Classification
@@
0.B.4 Open-Vocabulary Semantic Segmentation on Referring Datasets
We evaluate our model, which is fine-tuned TCL [4] with our method (TTD), on referring datasets [28, 42]. Unlike the text-level and tag-level segmentations, referring segmentation referring segmentation involves delineating entities described in free-form text. This task demands a deeper understanding of the textual context to accurately segment the corresponding entities in the image. As depicted in Tab. 11, our model surpasses existing approaches, showcasing its proficiency in comprehending textual descriptions and effectively translating them into precise image segmentations. This demonstrates the robustness and versatility of our method across diverse segmentation tasks, further validating its potential for real-world applications where understanding complex textual descriptions is paramount.
@@ 0.B.4 Open-Vocabulary Semantic Segmentation on Referring Datasets
0.B.5 Additional Refinement with SAM
@@
Appendix 0.C Generating Tag and Mask Annotations
@@
Appendix 0.D Additional Qualitative Results

### v2

@@ 4.1.1 Multi-Tag Selection
While image-tag scoring ($s_{\mathrm{image}}$) demonstrates high precision, it experiences very low recall due to the single tag bias, often leading to the selection of only one tag so if misses other image-related tags. Similarly, text-tag scoring ($s_{\mathrm{text}}$) and their combination show comparable issues. However, our approach significantly mitigates the single tag bias, showing high recall. Also, it exhibits high accuracy and mAP, which indicates the effectiveness of our scoring method. We further examine sensitivity to different thresholds and alternative scoring techniques that incorporate external data beyond the image-text pair in Sec. 0.D.1.
@@ 4.2 Results on Open-Vocabulary Semantic Segmentation
Tab. 4 contrasts the tag-level segmentation performance of studies employing only image-text pairs without additional annotations or external models akin to ours for fair comparison. For this reason, models [23, 11, 27, 13] that utilize additional annotations or external models (e.g., mask proposal networks) are also excluded. Additionally, we compared models that utilize NLP-based $\mathcal{L}$ in the table) for tag extraction. Implementing our fine-tuning technique results in an average increase in mIoUs of 8.5% for MaskCLIP and 3.5% for TCL. Furthermore, TCL enhanced with our method outperforms existing models on benchmarks such as Pascal VOC, Pascal Context, and COCO-Object. This indicates that our approach successfully augments model alignment in addition to alleviating single tag bias.
@@ 4.3 Ablation
Table 5: Ablation Studies.
(a) Effect of Loss Terms.
@@
(b) Effect of Tagging Method.
@@ 4.3.2 Effect of Tag Selection Method
In our approach, tag selection is conducted via pixel-tag scoring (Eq. 2) to extract pseudo-tags representing the image-text relationship. Alternatively, we examine NLTK and simple image-tag scoring (Eq. 1) for tag selection, as shown in Tab. 5(b). When employing NLTK to extract nouns from texts, due to the aforementioned issues (extraction of image-irrelevant tags and omission of image-relevant tags), the model struggles to generate the ideal text-image similarity map (i.e., the union of pseudo-tag driven maps). Consequently, the generated union map may include the activation of unrelated tags or omit activations necessary for the image-tag relationship, resulting in minor improvement. On the other hand, using image-tag scoring tends to extract a single tag, exacerbating single tag bias and resulting in a decline in both CaptionIoU and mIoU. In contrast, our tag selection method exhibits significantly enhanced performance in both tasks, indicating that the pseudo-tags extracted using our method effectively capture the image-text relationship.
@@ 5 Conclusion
Our research tackles the disproportionate alignment between images and text inherent in existing CLIP-based models, which we term as single tag bias. To address this issue, we introduce a novel fine-tuning method TTD that uses only image-text pairs without additional annotations or external models. We identify all ground-truth tags related to the image within the text and then generate an ideal image-text map that reflects all related tags for self-distillation. Our method effectively mitigates single tag bias, improves overall image-text alignment, and enhances performance in both image segmentation and multi-tag classification within the text. Our approach, following the TTD framework, enables the automatic generation of image-relevant tags, corresponding masks, and text masks, without reliance on external models beyond CLIP. This positions our method as a valuable tool for precise auto-labeling, with significant potential across various datasets consisting solely of images and texts.
Appendix 0.A Automatic Labeling via TTD
Table 6: Comparison of Our Method with Related Approaches.
Properties CoCu OVSeg TagAlign CaSED GroupViT MaskCLIP SimSeg TCL Ours
Automatic Annotations
(a) Generate tags from text
(b) Generate tag-level mask
(c) Generate text-level mask
Methodology
(d) Alleviate single tag bias
(e) No use of external models
(f) Learning pixel-tag alignment
(g) Follow model-agnostic manner
Figure 8: Automatic Labeling via TTD. The proposed TTD framework enables the automatic generation of (a) relevant tags from the input text, (b) corresponding masks for each tag, and (c) a composite/debiased mask representing the entire text.
Table 7: Performance of Automatic Labeling on the Train Set. Training CLIP variants (e.g., MaskCLIP/TCL) with our TTD improves (a) Multi-Tag Selection and (b) Text-level Semantic Segmentation. We assess performance based on the manually collected tag and mask annotations of the train set (CC3M and CC12M datasets) using the same metrics as Tabs. 2 and 3.
(a) Multi-Tag Selection.
Method Precision Recall F1 Accuracy
MaskCLIP 46.9 67.7 55.4 58.1
+ TTD 62.5 78.2 69.5 73.5 (+15.4)
TCL 78.0 83.0 80.4 84.4
+ TTD 80.7 82.1 81.4 85.5 (+1.1)
(b) Text-level Semantic Segmentation.
Method CaptionIoU mFPR mFNR
MaskCLIP 59.1 0.161 0.248
+ TTD 62.9 (+3.8) 0.172 0.199
TCL 71.0 0.147 0.144
+ TTD 72.2 (+1.2) 0.133 0.144
In this paper, we introduce TTD as a novel method capable of automatically labeling tag- and pixel-level annotations for image-text pairs from the training set, without the need for external pre-trained models beyond the baseline (CLIP), as illustrated in Tab. 6. Our approach autonomously selects image-relevant tags from the text (Fig. 8(a)) and generates a corresponding text mask (Fig. 8(c)), highlighting the image areas associated with the text based on tag masks (Fig. 8(b)). By eschewing external NLP or vision models in this process, we avoid biases stemming from their training.
Evaluation in Tab. 7 demonstrates the superior quality of tag/mask labels generated by TTD for the train set (CC3M and CC12M datasets). By mitigating single tag bias, our method outperforms previous approaches significantly. Our approach’s significance lies in its capacity to automatically generate rich annotations (tags/masks) across various training datasets (e.g., image captioning, medical), serving as a cost-efficient and precise auto-labeler, which is highly beneficial for CLIP-based applications. Thus, it facilitates learning pixel-tag alignment through these automatically generated labels.
Appendix 0.B Additional Analysis on Single Tag Bias
Figure 9: Distribution of Order for Word with Single Tag Bias. Word order within each sentence is normalized from 0 to 1. (e.g., the second word in a three-word caption has a normalized order of 0.5.)
We first identify the single tag bias, occurring in nearly $82\%$ of image-text pairs in the CC3M validation dataset (our proposed dataset). This frequency is determined using $s_{\mathrm{image}}$ (Eq. 1), where a single tag is selected despite the existence of multiple ground-truth tags, based on non-parametric thresholding introduced in Sec. 3.1.
Various factors, such as model overfitting and characteristics of the training data, may contribute to single tag bias. It might seem that bias occurs towards words describing elements deemed important in images (e.g., main object), often placed at the beginning of sentences in human-generated descriptions. However, an analysis of the biased tag positions in Fig. 9 indicates a scattered distribution throughout sentences, lacking significant correlation with specific positions. Investigating these underlying causes could be an intriguing avenue for future research.
Appendix 0.C Experimental Details
0.C.1 Implementation Details
@@
We conduct model training using 4 NVIDIA A100 GPUs, each having 80GB of memory. Implementation details are summarized in Tab. 8. Most settings, including image size, prompt template usage (for both tag and text input), and post-processing, align with the original setup of baseline models [44, 4]. Our batch size is set to 32, the maximum feasible size given our available GPU resources. Training data consists of the CC3M and CC12M datasets [34, 5], with additional sample pruning to remove noisy samples. We calculate cosine similarity between image and text embeddings for all pairs, sampling those with similarity values exceeding the mean plus the standard deviation. Ultimately, 1.8M pairs are selected from a total of 12.8M pairs, as shown in Tab. 9. Training time has been reduced approximately sevenfold, with a performance boost of about +0.6 mIoU, underscoring the efficacy of filtering out noisy samples.
@@ 0.C.2 Evaluation Details
Appendix 0.D Additional Experiments
@@
0.D.1 Effect of Tagging Methods
@@
0.D.2 Multi-Tag Selection
Here, we delve into an additional scoring method and analysis of multi-tag selection not covered in the main paper. We explore the performance of previously introduced scoring techniques ($s_{\mathrm{image}}$ (Eq. 1), $s_{\mathrm{text}}$ (Eq. 7)) across various thresholds in Sec. 0.C.2.
As mentioned in Sec. 1, each tag can be scored based on the proportion of image pixels similar to that tag ($s_{\mathrm{seg}}$). Specifically, as depicted in Fig. 11, we generate a similarity map between the image and each tag. Subsequently, for every pixel, we identify the most relevant tag with the highest similarity value, enabling us to score each tag based on the proportion of its relevant pixels. As illustrated in Tab. 10, $s_{\mathrm{seg}}$ demonstrates relatively high precision since it accentuates a single tag. However, due to the tendency of irrelevant tags to occupy more pixels than relevant ones, it encounters challenges in score ranking, resulting in a significantly low mAP (i.e., 57.9). In contrast, our scoring method ($s_{\mathrm{pixel}}$) effectively mitigates the single tag bias while maintaining correct ranking, resulting in a high F1 score and mAP.
Figure 12: Comparison with POS tagger and our TTD.
Table 12: Multi-Tag Selection Performance.
Method Encoding tags with templates P R F1 Acc
(a) NLTK (semantically meaningful tags) - 43.5 99.5 60.6 64.5
(b) TCL+TTD 87.3 75.4 80.9 90.1
(c) TCL+TTD 88.3 78.0 82.8 (+1.9) 91.0 (+0.9)
Existing tag selection methods rely solely on NLP modules to extract nouns from the given text. While it may seem intuitive to extract all semantically meaningful tags beyond nouns, many of them may not be relevant to the image content, as shown in Fig. 12 and Tab. 12(a). For example, even in cases where a woman is absent from the image, the POS tagger still identifies it as an image-related tag. (second image in Fig. 12) In contrast, our TTD accurately extracts image-relevant tags, demonstrating its superiority over the simple extraction methods using POS tagger (Tab. 12(c)).
Furthermore, we examine the influence of text templates on encoding tags and text inputs. Our findings in Tab. 12 demonstrate that template usage improves the performance of multi-tag selection. This aligns with the observation that CLIP typically excels at encoding complete sentences or descriptions rather than individual tags.
@@ 0.D.2 Multi-Tag Selection
ADDS [40] ViT-B/32 36.56
MKT [15] ViT-B/16 37.6
DualCoOp [36] ResNet-50 43.6
TaI-DPT [14] ResNet-50 44.99
CLIP [31] ResNet-50 35.58
CLIP Surgery [22] ResNet-50 39.99
CLIP Surgery [22] ViT-B/16 40.53
Ours (TCL [4]+TTD) ViT-B/16 42.63
0.D.3 Open-Vocabulary Multi-label Classification
@@
0.D.4 Open-Vocabulary Semantic Segmentation on Referring Datasets
We evaluate our model, which is fine-tuned TCL [4] with our method (TTD), on referring datasets [28, 42]. Unlike the text-level and tag-level segmentations, referring segmentation involves delineating entities described in free-form text. This task demands a deeper understanding of the textual context to accurately segment the corresponding entities in the image. As depicted in Tab. 14, our model surpasses existing approaches, showcasing its proficiency in comprehending textual descriptions and effectively translating them into precise image segmentations. This demonstrates the robustness and versatility of our method across diverse segmentation tasks, further validating its potential for real-world applications where understanding complex textual descriptions is paramount.
@@ 0.D.4 Open-Vocabulary Semantic Segmentation on Referring Datasets
0.D.5 Additional Refinement with SAM
@@
Appendix 0.E Generating Tag and Mask Annotations
@@
Appendix 0.F Additional Qualitative Results